# Emergent spin order and steady-state superradiance in one-dimensional baths

### v1

@@ IV.1 Thresholds and optimal intensity
Figure 5: Thresholds for superradiant light emission and scaling of maximum emission rate. (a) Emission rate and (b) excited-state population in the steady state as a function of pumping strength. Insets: zoom on the first threshold. (c) Maximum intensity vs atom number. In all plots, solid lines show the analytical expressions [from Ref. [Meiser10] for the single-mode cavity and from Eqs. (33) for the ring cavity], points show the numerical results with TWA, and the dashed line in (c) indicates the estimate of $R_{\text{max}}$ for the waveguide described in the main text. Parameters: $(N,\Gamma^{\prime},kd)=(1024,2\Gamma_{\rm 1D},2\pi/3)$.
In a cavity, the lower and upper thresholds for superradiant lasing are known to be $w^{(l)}=\Gamma^{\prime}$ and $w^{(u)}=N\Gamma_{\rm 1D}$, respectively, with a maximum emission rate scaling as $R_{\text{max}}=N^{2}\Gamma_{\rm 1D}/8$ [Meiser09, Meiser10]. The thresholds are intuitively clear since incoherent pumping must exceed the independent decay rate to generate positive inversion and enable synchronization, while excessively strong pumping destroys correlations seeded by collective decay. Since $N\Gamma_{\rm 1D}$ sets the maximum collective decay timescale, for $w>w^{(u)}$ correlations are washed out faster than they can form.
We find analytical expressions for the ring-cavity thresholds for superradiant emission by exploiting the superspin method introduced in Ref. [Lee25] and described in Sec. II. Analyzing the Heisenberg evolution of the collective operators in Eq. (14) and performing a second-order cumulant expansion analogous to the single-mode case (a summary of this calculation is given in Appendix A), we find that the ring-cavity steady state also exhibits threshold behavior. In particular, the steady-state total intensity and excited–state population $P_{e}\equiv\sum_{n}\langle\hat{\sigma}_{ee}^{n}\rangle$ read
@@
Figures 5(a,b) show good agreement between these expressions and the values obtained with TWA. Numerical results for non-commensurate configurations with $d\neq m\lambda_{0}/2p$ follow the same behavior, indicating that the steady-state values of these observables are largely insensitive to the spatial arrangement of the atoms. Equations (33) predict sharp changes in population and the onset and quenching of steady-state collective emission at a lower threshold $w^{(\ell)}=\Gamma^{\prime}$ and an upper threshold $w^{(u)}=N\Gamma_{\text{1D}}/2$. The insets of Fig. 5(a,b) zoom into the behavior around the first threshold. Below $w^{(\ell)}$, TWA yields a small unphysical negative decay rate, consistent with the fact that the steady state in this regime is known to be dark and highly entangled for a single-mode cavity [Shankar21], a scenario expected to carry over to 1D reservoirs. Despite this limitation, TWA correctly captures the location of $w^{(\ell)}$. Moreover, Eq. (33a) predicts a maximum emission rate at $w=N\Gamma_{\text{1D}}/4$, with $R_{\text{max}}=N^{2}\Gamma_{\text{1D}}/16$, in agreement with numerical results as shown in Fig. 5(c). As for a single-mode cavity, the upper threshold reflects the requirement that the pump strength remain below the maximum dissipation rate of the system [Meiser10, Mok24], which also sets the fastest synchronization timescale in Eq. (23).
@@
For the waveguide, we obtain the thresholds numerically. These results are in good agreement with a simple back-of-the-envelope analytical calculation, described below. We find $w^{(\ell)}\sim\Gamma^{\prime}$ for all array spacings, but the exact position of the upper threshold depends on lattice constant. Away from the “mirror” configuration, we estimate the upper threshold from Eqs. (29). Because the feedback is position-dependent, each atom has a different maximum pumping at which it can remain synchronized. For instance, atom 1 experiences feedback from (N1)(N-1) atoms driving it toward the left order, so the maximum pumping strength for synchronization is $w^{(u)}(n=1)\sim(N-1)\Gamma_{\rm 1D}/2$. More generally, an atom in the first half of the array will synchronize only if $w<w^{(u)}(n)\sim\left(\tfrac{N}{2}-n+1\right)\Gamma_{\rm 1D}$. Averaging over the array then yields $\bar{w}^{(u)}\sim N\Gamma_{\rm 1D}/4$, and by similar reasoning, an optimal pumping $\bar{w}^{(opt)}\sim N\Gamma_{\rm 1D}/8$. Counting atoms with $w^{(u)}(n)<\bar{w}^{(opt)}$ that synchronize to either order at the optimal pump (denoted $N_{\text{sync}}$) and using the known single-mode cavity scaling, we estimate the waveguide intensity at optimal pumping to be $R_{\text{max}}\sim 2\left(N_{\text{sync}}^{2}\Gamma_{\rm 1D}/8\right)=9N^{2}\Gamma_{\rm 1D}/256$. These estimates qualitatively agree with the numerical results in Fig. 5. In contrast to transient superradiance, where coherent interactions can be often neglected when analyzing the superradiant burst [Masson22, Cardenas23], the coherent interactions intrinsic to the waveguide strongly modifies the properties of the radiation emitted in the steady state.
@@ IV.2 Spectral properties
Figure 6: Linewidth of the emitted field. (a) Linewidth vs. pumping strength for $N=50$ atoms for the single-mode cavity, ring cavity and waveguide models. The solid lines show exact calculations using $SU(4)$ methods and the dots are obtained by means of the TWA. (b) Scaling of the linewidth vs NN. Continuous lines are guides to the eye, dashed line is exact, computed via $SU(4)$ methods. In both plots $(kd,\Gamma^{\prime})=(2\pi/3,2\Gamma_{\rm 1D})$.
In a single-mode cavity, the emitted light is coherent and there is a superradiant “lasing” regime if the pumping is between thresholds [Meiser09]. This means that the order emerging from synchronization in a single-mode cavity is robust against fluctuations, allowing steady-state correlations to persist for long times [Meiser09]. The two-time correlation of any pair of atoms decays exponentially with the time delay $\tau$, i.e.,
@@
with a coherence time $t_{\text{coh}}$ that increases with $N$ [Meiser10]. The enhanced lifetime is directly reflected in the spectrum of the light emitted by the cavity, given for large atom number by
$S_{\hat{E}^{+}_{C}}(\omega)\simeq N(N-1)\,\text{Re}\!\left[\int_{0}^{\infty}d\tau\,e^{-{\rm i}\omega\tau}\langle\hat{\sigma}_{-}^{1}(t_{ss}+\tau)\hat{\sigma}_{+}^{2}(t_{ss})\rangle\right],$ (35)
which exhibits a linewidth much narrower than that of independent atoms, $(\Delta\nu)_{\text{indep}}=w+\Gamma^{\prime}+\Gamma_{\rm 1D}$. In the large-$N$ limit, the linewidth approaches $(\Delta\nu)_{\text{min}}\sim\Gamma_{\rm 1D}$, so correlations can persist for a maximum time $t_{\text{coh}}\sim 1/\Gamma_{\rm 1D}$ even when parasitic decay dominates over cavity coupling, $\Gamma^{\prime}\gg\Gamma_{\rm 1D}$ [Meiser09]. The residual broadening is primarily due to phase fluctuations of the emitted field [Xu16]: as shown in Fig. 2, the amplitude is fixed while the phase remains undetermined. Reduced phase noise enables such light to be used to lock a laser, or as an active frequency reference.
As discussed in Sec. III, synchronization beyond the single-cavity similarly generates phase correlations across the atomic array, but it remains unclear whether these emerging orders are also capable to radiate light with phase robustness that is strengthened as $N$ grows. To address this question, we analyze steady-state correlations of the right- and left-emitted fields. Figure 6 summarizes the linewidth behavior as a function of pump strength and atom number for the three spin models. For moderate $N$, the linewidth exhibits a non-monotonic dependence on the pump rate within the synchronization window [Fig. 6(a)], revealing an optimal pump rate that minimizes the linewidth. Linewidths are computed using TWA [Guardiola25] and validated against exact $SU(4)$ calculations for the cavity and ring cavity, which reproduce the same qualitative trends.
Figure 6(b) shows the minimum linewidth as a function of atom number. Both the spectral linewidth of $\hat{E}^{+}_{C}$ in the single-mode cavity and that of $\hat{E}^{+}_{RC,R/L}$ in the ring cavity exhibit a seemingly downward trend with increasing $N$. In the waveguide, we calculate the linewidth of the right (left) field at an observation point to the right (left) of all atoms. Our numerical estimations do not conclusively show line narrowing, as the results are too noisy (see Appendix C for more details). This ambiguity may stem from numerical limitations; for example, the narrowing may only appear at atom numbers beyond our simulation capabilities, or the noise inherent to the TWA and our signal processing methods may obscure the physical trend. Alternatively, this lack of narrowing may be physical, indicating that the spin order in the waveguide is not robust against quantum fluctuations.
@@ V SUMMARY AND OUTLOOK
We have studied steady-state superradiance of incoherently pumped atoms coupled to 1D reservoirs, focusing on two setups: a ring cavity, where atoms collectively decay into left- and right-propagating modes, and a bidirectional waveguide, where coherent exchange processes accompany collective decay. In the ring cavity, individual realizations lock to one of two macroscopic orders that maximize emission into a single direction, whereas in the waveguide coherent interactions enforce coexistence of left and right order, with edge atoms synchronizing to opposite modes. This self-organization shapes the emitted light. For the ring cavity we derived analytical thresholds and optimal intensities, and observed a downward trend in the linewidth of the emitted light. For the waveguide we obtained numerical estimates of thresholds and intensities, but no clear line narrowing was observable within the sytem sizes available to our numerical methods.
@@
Our results suggest several directions for future work. On the experimental side, they motivate studies of steady-state superradiance in ring resonators, bow-tie cavities, and nanofibers. On the theory side, a more exhaustive analysis of the linewidth in the waveguide setup is needed, for instance through cumulant expansions. Beyond this, it would be valuable to characterize shot-to-shot fluctuations in the ordering through semiclassical simulations of individual trajectories [Yu24, Zhang25] to clarify the stability of the phase-ordered state and the origin, or possible absence, of linewidth narrowing. Finally, it would be interesting to determine whether any correlator becomes increasingly robust to fluctuations as $N$ grows.
@@
Finally, a natural extension of this work is to explore the prospects for superradiant lasing in free space, as originally proposed by Dicke [Dicke64]. In particular, it would be interesting to see whether collective decay in the absence of coherent interactions generically leads to basins of attraction associated with the brightest collective modes, as in the ring cavity, and whether the coherent interactions generate crossovers between competing directional orders, as observed in the waveguide. Recent work [Bychek25] has shown that a few atoms in a 1D array can emit light with a linewidth that converges to the natural single-atom value by limiting pump-induced broadening and driving only a subset of atoms. In practice, however, 1D atomic arrays with ultranarrow transitions may not be suitable to generate light for atomic clock interrogation, since their maximum collective decay rate scales only linearly with $N$ [Mok24], preventing the realization of superradiant lasers (as the emission is too weak even for large systems). Nevertheless, it would be worth investigating whether selectively pumping only part of a higher-dimensional array could overcome this limitation and enable superradiant lasing in free-space geometries.
@@
Acknowledgments – We are grateful to Joseph T. Lee and Ricardo Gutierrez-Jauregui for discussions. We acknowledge support by the National Science Foundation through the CAREER Award (No. 2047380), the Air Force Office of Scientific Research through their Young Investigator Program (grant No. 21RT0751), the David and Lucile Packard Foundation, and the Chu Family Foundation.
@@ Appendix A LARGE SPIN LIMIT AND THRESHOLDS
Here $\hat{H}$ is the Hamiltonian defined in Eq. (2a), and $\mathcal{F}\left[\hat{\mathcal{O}}\right$] is quantum Langevin noise, which averages to zero and can be neglected when computing expectation values of $\hat{\mathcal{O}}$ [Rubies23].
As mentioned in the main text, for atoms coupled to a ring cavity, spacings satisfying $d=m\lambda_{0}/2p$, with $m,p\in\mathbb{Z}$ exhibit partial permutational symmetry. We therefore represent each permutationally symmetric group $\mathcal{S}_{\alpha}$ ($\alpha=1,\dots,p$) by macroscopic angular-momentum–like operators [Lee25]
@@
where $c_{n,\alpha}=\lfloor(n-1)/\alpha\rfloor+1$ counts how many wavelengths atom $n$ is away from the first atom in its symmetry group, so that the factor $(-1)^{m(c_{n,\alpha}+1)}$ adds contributions within superspin $\mathcal{S}_{\alpha}$ in phase. Notice that because of the $1/N_{\alpha}$ normalization these operators do not close a $SU(2)$ algebra.
Evaluating Eq. (36) for these operators and applying a second-order cumulant expansion [Plankensteiner22, Rubies23], the dynamics of $p$ incoherently driven superspins reads
@@
For $p=1$ and large $N$, Eqs. (39) reduce to the superradiant laser equations of Ref. [Meiser10], i.e.,

### v2

@@ IV.1 Thresholds and optimal intensity
Figure 5: Thresholds for superradiant light emission and scaling of maximum emission rate. (a) Emission rate and (b) excited-state population in the steady state as a function of pumping strength. Insets: zoom on the first threshold. (c) Maximum intensity vs atom number. In all plots, solid lines show the analytical expressions [from Ref. [41] for the single-mode cavity and from Eqs. (33) for the ring cavity], points show the numerical results with TWA, and the dashed line in (c) indicates the estimate of $R_{\text{max}}$ for the waveguide described in the main text. Parameters: $(N,\Gamma^{\prime},kd)=(1024,2\Gamma_{\rm 1D},2\pi/3)$.
In a cavity, the lower and upper thresholds for superradiant lasing are known to be $w^{(l)}=\Gamma^{\prime}$ and $w^{(u)}=N\Gamma_{\rm 1D}$, respectively, with a maximum emission rate scaling as $R_{\text{max}}=N^{2}\Gamma_{\rm 1D}/8$ [42, 41]. The thresholds are intuitively clear since incoherent pumping must exceed the independent decay rate to generate positive inversion and enable synchronization, while excessively strong pumping destroys correlations seeded by collective decay. Since $N\Gamma_{\rm 1D}$ sets the maximum collective decay timescale, for $w>w^{(u)}$ correlations are washed out faster than they can form.
We find analytical expressions for the ring-cavity thresholds for superradiant emission by exploiting the superspin method introduced in Ref. [34] and described in Sec. II. Analyzing the Heisenberg evolution of the collective operators in Eq. (14) and performing a second-order cumulant expansion analogous to the single-mode case (a summary of this calculation is given in Appendix A), we find that the ring-cavity steady state also exhibits threshold behavior. In particular, the steady-state total intensity and excited–state population $P_{e}\equiv\sum_{n}\langle\hat{\sigma}_{ee}^{n}\rangle$ read
@@
Figures 5(a,b) show good agreement between these expressions and the values obtained with TWA. Numerical results for non-commensurate configurations with $d\neq m\lambda_{0}/2p$ follow the same behavior, indicating that the steady-state values of these observables are largely insensitive to the spatial arrangement of the atoms. Equations (33) predict sharp changes in population and the onset and quenching of steady-state collective emission at a lower threshold $w^{(\ell)}=\Gamma^{\prime}$ and an upper threshold $w^{(u)}=N\Gamma_{\text{1D}}/2$. The insets of Fig. 5(a,b) zoom into the behavior around the first threshold. Below $w^{(\ell)}$, TWA yields a small unphysical negative decay rate, consistent with the fact that the steady state in this regime is known to be dark and highly entangled for a single-mode cavity [56], a scenario expected to carry over to 1D reservoirs. Despite this limitation, TWA correctly captures the location of $w^{(\ell)}$. Moreover, Eq. (33a) predicts a maximum emission rate at $w=N\Gamma_{\text{1D}}/4$, with $R_{\text{max}}=N^{2}\Gamma_{\text{1D}}/16$, in agreement with numerical results as shown in Fig. 5(c). As for a single-mode cavity, the upper threshold reflects the requirement that the pump strength remain below the maximum dissipation rate of the system [41, 45], which also sets the fastest synchronization timescale in Eq. (23).
@@
For the waveguide, we obtain the thresholds numerically. These results are in good agreement with a simple back-of-the-envelope analytical calculation, described below. We find $w^{(\ell)}\sim\Gamma^{\prime}$ for all array spacings, but the exact position of the upper threshold depends on lattice constant. Away from the “mirror” configuration, we estimate the upper threshold from Eqs. (29). Because the feedback is position-dependent, each atom has a different maximum pumping at which it can remain synchronized. For instance, atom 1 experiences feedback from (N1)(N-1) atoms driving it toward the left order, so the maximum pumping strength for synchronization is $w^{(u)}(n=1)\sim(N/2-1)\Gamma_{\rm 1D}$. More generally, an atom in the first half of the array will synchronize only if $w<w^{(u)}(n)\sim\left(N/2-n\right)\Gamma_{\rm 1D}$. Averaging over the array then yields $\bar{w}^{(u)}\sim N\Gamma_{\rm 1D}/4$, and by similar reasoning, an optimal pumping $\bar{w}^{(opt)}\sim N\Gamma_{\rm 1D}/8$. Counting atoms with $w^{(u)}(n)<\bar{w}^{(opt)}$ that synchronize to either order at the optimal pump (denoted $N_{\text{sync}}$) and using the known single-mode cavity scaling, we estimate the waveguide intensity at optimal pumping to be $R_{\text{max}}\sim 2\left(N_{\text{sync}}^{2}\Gamma_{\rm 1D}/8\right)=9N^{2}\Gamma_{\rm 1D}/256$. These estimates qualitatively agree with the numerical results in Fig. 5. In contrast to transient superradiance, where coherent interactions can be often neglected when analyzing the superradiant burst [39, 11], the coherent interactions intrinsic to the waveguide strongly modifies the properties of the radiation emitted in the steady state.
@@ IV.2 Spectral properties
Figure 6: Linewidth of the emitted field. (a) Linewidth vs. pumping strength for $N=50$ atoms for the single-mode cavity, ring cavity and waveguide models. (b) Scaling of the linewidth vs NN. In both panels, points are obtained using the TWA, dashed lines show exact results based on the $SU(4)$ formalism, and solid lines serve as guides to the eye. For the ring cavity and waveguide, $(kd,\Gamma^{\prime})=(2\pi/3,2\Gamma_{\rm 1D})$.
In a single-mode cavity, the emitted light is coherent and there is a superradiant “lasing” regime if the pumping is between thresholds [42]. This means that the order emerging from synchronization in a single-mode cavity is robust against fluctuations, allowing steady-state correlations to persist for long times [42]. The two-time correlation of any pair of atoms decays exponentially with the time delay $\tau$, i.e.,
@@
with a coherence time $t_{\text{coh}}$ that increases with $N$ [41]. The enhanced lifetime is directly reflected in the spectrum of the light emitted by the cavity, given for large atom number by
$S_{\hat{E}^{+}_{C}}(\omega)\simeq N(N-1)\,\text{Re}\!\left[\int_{0}^{\infty}d\tau\,e^{-{\rm i}\omega\tau}\langle\hat{\sigma}_{-}^{1}(t_{ss}+\tau)\hat{\sigma}_{+}^{2}(t_{ss})\rangle\right],$
which exhibits a linewidth much narrower than that of independent atoms, $(\Delta\nu)_{\text{indep}}=w+\Gamma^{\prime}+\Gamma_{\rm 1D}$. In the large-$N$ limit, the linewidth approaches $(\Delta\nu)_{\text{min}}\sim\Gamma_{\rm 1D}$, so correlations can persist for a maximum time $t_{\text{coh}}\sim 1/\Gamma_{\rm 1D}$ even when parasitic decay dominates over cavity coupling, $\Gamma^{\prime}\gg\Gamma_{\rm 1D}$ [42]. The residual broadening is primarily due to phase fluctuations of the emitted field [68]: as shown in Fig. 2, the amplitude is fixed while the phase remains undetermined. Reduced phase noise enables such light to be used to lock a laser, or as an active frequency reference.
As discussed in Sec. III, synchronization beyond the single-cavity similarly generates phase correlations across the atomic array, but it remains unclear whether these emerging orders are also capable to radiate light with phase robustness that is strengthened as $N$ grows. To address this question, we analyze steady-state correlations of the right- and left-emitted fields. Figure 6 summarizes the linewidth behavior as a function of pump strength and atom number for the three spin models. For moderate $N$, the linewidth exhibits a non-monotonic dependence on the pump rate within the synchronization window [Fig. 6(a)], revealing an optimal pump rate that minimizes the linewidth. Linewidths are computed using TWA [26] and validated against exact $SU(4)$ calculations for the cavity, which reproduce the same qualitative trends.
Figure 6(b) shows the minimum linewidth as a function of atom number. Both the spectral linewidth of $\hat{E}^{+}_{C}$ in the single-mode cavity and that of $\hat{E}^{+}_{RC,R/L}$ in the ring cavity exhibit a seemingly downward trend with increasing $N$. In the waveguide, we calculate the linewidth of the right (left) field at an observation point to the right (left) of all atoms. After accounting for the estimation uncertainty (see Appendix C for details), our numerical results do not conclusively demonstrate linewidth narrowing. This inconclusiveness may arise from numerical limitations. For instance, comparison with exact linewidths in the cavity case, obtained using an $SU(4)$ representation of the dynamics [67], shows that the TWA captures the true linewidth narrowing only qualitatively. In addition, narrowing may only emerge at atom numbers beyond our current simulation capabilities. Alternatively, the absence of clear narrowing may be a genuine physical effect, indicating that the spin order in the waveguide is not robust enough to eliminate noise from the parasitic decay.
@@ V SUMMARY AND OUTLOOK
We have studied steady-state superradiance of incoherently pumped atoms coupled to 1D reservoirs, focusing on two setups: a ring cavity, where atoms collectively decay into left- and right-propagating modes, and a bidirectional waveguide, where coherent exchange processes accompany collective decay. In the ring cavity, individual realizations lock to one of two macroscopic orders that maximize emission into a single direction, whereas in the waveguide coherent interactions enforce coexistence of left and right order, with edge atoms synchronizing to opposite modes. This self-organization shapes the emitted light. For the ring cavity we derived analytical thresholds and optimal intensities, and observed a downward trend in the linewidth of the emitted light. For the waveguide we obtained numerical estimates of thresholds and intensities, but no clear line narrowing was observable within the system sizes available to our numerical methods.
@@
Our results suggest several directions for future work. On the experimental side, they motivate studies of steady-state superradiance in ring resonators, bow-tie cavities, and nanofibers. On the theory side, a more exhaustive analysis of the linewidth in the waveguide setup is needed, for instance through cumulant expansions. Beyond this, it would be valuable to characterize shot-to-shot fluctuations in the ordering through semiclassical simulations of individual trajectories [70, 72] to clarify the stability of the phase-ordered state and the origin, or possible absence, of linewidth narrowing. It would be interesting to determine whether any correlator becomes increasingly robust to fluctuations as $N$ grows.
@@
Finally, a natural extension of this work is to explore the prospects for superradiant lasing in free space, as originally proposed by Dicke [19]. In particular, it would be interesting to see whether collective decay in the absence of coherent interactions generically leads to basins of attraction associated with the brightest collective modes, as in the ring cavity, and whether the coherent interactions generate crossovers between competing directional orders, as observed in the waveguide. Recent work [9] has shown that a few atoms in a 1D array can emit light with a linewidth that converges to the natural single-atom value by limiting pump-induced broadening and driving only a subset of atoms. In practice, however, 1D atomic arrays with ultranarrow transitions may not be suitable to generate light for atomic clock interrogation, since their maximum collective decay rate scales only linearly with $N$ [45], preventing the realization of superradiant lasers (as the emission is too weak even for large systems). Nevertheless, it would be worth investigating whether selectively pumping only part of a higher-dimensional array could overcome this limitation and enable superradiant lasing in free-space geometries.
@@
Acknowledgments – We are grateful to Joseph T. Lee, Ricardo Gutierrez-Jauregui, Xueyue Zhang and Tingyao Wang for discussions. We acknowledge support by the National Science Foundation through the CAREER Award (No. 2047380), the Air Force Office of Scientific Research through their Young Investigator Program (grant No. 21RT0751), the David and Lucile Packard Foundation, and the Chu Family Foundation.
@@ Appendix A LARGE SPIN LIMIT AND THRESHOLDS
Here $\hat{H}$ is the Hamiltonian defined in Eq. (2a), and $\mathcal{F}\left[\hat{\mathcal{O}}\right$] is quantum Langevin noise, which averages to zero and can be neglected when computing expectation values of $\hat{\mathcal{O}}$ [52].
As mentioned in the main text, for atoms coupled to a ring cavity, spacings satisfying $d=m\lambda_{0}/2p$, with $m,p\in\mathbb{Z}$ exhibit partial permutational symmetry. We therefore represent each permutationally symmetric group $\mathcal{S}_{\alpha}$ ($\alpha=1,\dots,p$) by macroscopic angular-momentum–like operators [34]
@@
where $c_{n,\alpha}=\lfloor(n-1)/\alpha\rfloor+1$ counts how many wavelengths atom $n$ is away from the first atom in its symmetry group, so that the factor $(-1)^{m(c_{n,\alpha}+1)}$ adds contributions within superspin $\mathcal{S}_{\alpha}$ in phase. Notice that because of the $1/N_{\alpha}$ normalization these operators do not form a $SU(2)$ algebra.
Evaluating Eq. (36) for these operators and applying a second-order cumulant expansion [48, 52], the dynamics of $p$ incoherently driven superspins reads
@@
For $p=1$ and large $N$, Eqs. (39) reduce to the superradiant laser equations of Ref. [41], i.e.,